# The coherent structure of the energy cascade in isotropic turbulence

### v3

@@ Introduction
There is general agreement from the community that the answer to the first question is yes. The net transfer of energy from large to small scales is mainly accomplished by interactions among flow motions of similar size. The conclusion is supported by evidence from diverse approaches, such as scaling analysis [12, 13, 14, 15, 16, 17, 18, 19, 20], triadic interactions in Fourier space [21, 22], time correlations [23, 24], and information-theoretic causality [25], to name a few.
The degree of consensus is lower regarding the second question. Several theories have been proposed to describe the physical mechanism(s) that enable the transfer of energy from larger to smaller scales throughout the inertial range. One of the first mechanisms proposed is based on the concept of vortex stretching. In this scenario, vorticity is stretched by the strain-rate either at the same scale or at a larger scale [26, 27, 28, 29, 30, 31, 32, 33]. Some authors have further proposed that the mechanistic details of the process can be explained by successive reconnections of anti-parallel vortex tubes [34, 35, 36, 37], vortex reconnection of two long, straight anti-parallel vortex tubes with localized bumps [38], and the presence of helictical instabilities in characteristic of vortex rings [39, 40]. These viewpoints, while not dynamically equivalent, are still compatible with the vortex-stretching driven energy cascade.
The main competing theory to the vortex stretching mechanism is the self-amplification of the rate-of-strain either by same-scale interactions or by the amplification from larger scale strain-rate [41, 42, 43, 44, 45]. In both scenarios, the strain-rate self-amplification stands as the key contributor to the transfer of energy among scales, whereas vortex stretching is merely the effect (rather than the cause) of the energy cascade. Given the kinematic relationship between vortex stretching and strain-rate self-amplification [46, 47], some authors have argued that both mechanisms play a relevant role in the dynamics of the energy cascade [48, 49].
@@ Methods
We use direct numerical simulations of isotropic turbulence from Ref. [23]. The numerical setup corresponds to isotropically forced turbulence within a triply periodic domain at $\mathrm{Re}_{\lambda}=146,236$, and $384$, where $\mathrm{Re}_{\lambda}$ is the Reynolds number based on the Taylor microscale. The simulations are labeled as HIT1 (for $\mathrm{Re}_{\lambda}=146$), HIT2 (for $\mathrm{Re}_{\lambda}=236$), and HIT3 (for $\mathrm{Re}_{\lambda}=384$). Table 1 contains the numerical details of the simulations in terms of grid resolution, domain size, and number of snapshots. The reader is referred to Ref. [23] for additional information about the numerical setup.
@@
To investigate the structure of the energy transfer, the velocity field is decomposed into large and small scales using a low-pass Gaussian filter such that $u_{i}=\widetilde{u}_{i}+u^{\prime}_{i}$ for $i=1,2,3$, where ${u}_{i}$ represents the $i$-th component of the instantaneous velocity, $\widetilde{u}_{i}$ is the filtered velocity, and $u^{\prime}_{i}$ corresponds to the remaining fluctuating velocity component. The filter is isotropic and given in Fourier space by the Gaussian kernel $\hat{G}(\mathbf{\kappa})=\exp[-(r\kappa)^{2}/24]$, where $\kappa=\boldsymbol{|\kappa|}$ is the magnitude of the wavenumber vector ($\boldsymbol{\kappa}$), and $r$ is the filter width. Four filter widths are investigated: $r/\eta=20,30,40,$ and $60$, where $\eta=\left(\nu^{3}/\bar{\varepsilon}\right)^{1/4}$ is the Kolmogorov length-scale, $\nu$ is the kinematic viscosity, and $\bar{\epsilon}$ is the space-time averaged dissipation of turbulent kinetic energy. Figure 1 compares one instant of the unfiltered velocity and the filtered velocity for $r=30\eta$. The values of $r$ were chosen to span across the inertial range of the turbulence cascade [23]. The kinetic energy spectra $E(\kappa)$ for the three Reynolds numbers are shown in figure 2(a). As expected, the viscous and inertial range of $E(\kappa)$ collapse in Kolmogorov units, defined by $\eta$ and the Kolmogorov velocity-scale $u_{\eta}=(\nu\bar{\epsilon})^{1/4}$. The compensated kinetic energy spectra $\kappa^{5/3}E(\kappa)$ for HIT3 is included in figure 2(b) for the unfiltered case and for $r/\eta=20,30,40$ and $60$. The latter shows that the filtered widths lie within the inertial range of the moderate Reynolds numbers considered.
@@
The instantaneous inter-scale kinetic energy transfer at one instance is visualized in figure 3(a). Our primary efforts are directed toward analyzing the coherent structure of the flow in the vicinity of intense forward cascade events. Regions of strong positive $\Pi$ are isolated by thresholding the field with the parameter $\alpha>0$, such that $\Pi>\alpha\Pi_{\text{rms}}$, where $\Pi_{\text{rms}}$ represents the standard deviation of $\Pi$ over all the dataset. The value of $\alpha=1$ is selected following previous work [54]. We show in the Supplementary Information that the conclusions drawn in this manuscript also hold for other values of $\alpha$. Then, individual three-dimensional $\Pi$-structures are defined as contiguous regions in space satisfying $\Pi>\Pi_{\text{rms}}$. These structures will be utilized as markers to conditionally average the flow around them. The $\Pi$-structures occupy less than 20% of the total fluid volume but contribute to 70% of the total kinetic energy transfer. Figure 3(b) visualizes the regions associated with forward (red, $\Pi>\Pi_{\text{rms}}$) and backward (blue, $\Pi<-\Pi_{\text{rms}}$) intense cascade events for a given snapshot. Note that we use $\Pi$ to designate the scalar field of inter-scale kinetic energy transfer, which is a function of both space and time, whereas the term $\Pi$-structure refers to the individual spatial structures defined by contiguous regions in space that satisfy $\Pi>\Pi_{\text{rms}}$ at a given time. Forward cascade events dominate over their backward counterparts, as clearly appreciated in figure 3(b), consistent with previous results in the literature [57, 58, 59]. We will focus solely on the forward energy cascade, while we do not address inverse cascading events. This decision is primarily driven by two factors. Firstly, the number of intense backward cascade events is orders of magnitude smaller than the forward events, as illustrated in figure 3(b). Secondly, whereas the importance of the forward energy cascade is widely recognized in terms of dynamics and reduced-order modeling of the flow, the same consensus is not established for the inverse energy cascade in three-dimensional turbulence. Recent works have suggested that the inverse energy cascade may play a minor role in the flow dynamics or even lack physical significance [45, 55, 25].
@@
The typical shape of the $\Pi$-structures is quantified via their fractal dimension. We employed the box-counting method to calculate the Minkowski-Bouligand dimension of each $\Pi$-structure [60, 61]. For each object, the computational domain is divided into cubes with a side length of $l$, and the number $N_{b}(l)$ of boxes containing at least one point of the object is counted. In the case of a pure fractal set with dimension $D$, the number of boxes would follow a power law $N_{b}(l)\sim l^{-D}$. In practice, this relationship only holds within a restricted range of scales, bounded by a large-scale and a small-scale cutoff. We define the local fractal dimension as
@@
The fractal dimension of the object can be taken along the range over which the slope of $D_{l}(l)$ is approximately constant.
Figure 4(d) shows the PDF of the fractal dimensions for the $\Pi$-structures for different filter widths and Reλ. Space-filling spheroidal shapes would exhibit fractal dimensions close to $D\approx 3$, whereas sheet-like or filament-like structures would have $D\approx 2$ or $D\approx 1$, respectively. The PDF of $D$ reveals a bell-shaped distribution centered around $2.45$, suggesting that the objects have an intermediate form between sheet-like and strictly spheroidal shapes. Therefore, although some $\Pi$-structures may exhibit complex shapes, most often they resemble elongated ellipsoids with arbitrary orientation. In summary, the results from this section show that $\Pi$-structures are geometrically self-similar across scales within the inertial range and for the three Reλ considered. This finding provides justification for the conditional averaging procedure introduced in the subsequent section.
@@
    First, individual $\Pi$-structures are identified using the thresholding approach discussed in §Identification of intense energy transfer events. Examples of three individual $\Pi$-structures are shown in figure 5. A characteristic length scale is assigned to each $\Pi$-structure, defined by the diagonal of their bounding box, $\sqrt{l_{x}^{2}+l_{y}^{2}+l_{z}^{2}}$. Our interest lies in studying structures that are of a size similar to the filter width. $\Pi$-structures larger than half the length of the entire computational domain are excluded from the analysis. Similarly, structures smaller than $3^{3}$ grid points are also disregarded. The number of valid $\Pi$-structures considered for the analysis is roughly on the order of $\mathcal{O}(10^{4})$ for each Reλ and filter width. Given the large number of samples, the statistical uncertainty of the results were found to be small and more details are discussed in the Supplementary Information.
@@
    Each individual $\Pi$-structure is assigned a local frame of reference, with the origin positioned at its center of mass. This is the same frame of reference used in §Geometric properties of $\Pi$-structures to compute the characteristic lengths of the $\Pi$-structures. The axes directions align with the principal axes of inertia of the structure, determined by the eigenvectors of the moment of inertia tensor relative to the center of mass. A rotation matrix, constructed from these eigenvectors, is used to rotate the flow field surrounding each structure. Figure 5 illustrates the rotated bounding boxes for three $\Pi$-structures. The principal axes of inertia introduced above does not inherently define the positive or negative directions of the axes. To resolve this ambiguity, we set the direction of the axes such that the highest enstrophy $\widetilde{\omega}_{i}\widetilde{\omega}_{i}$, averaged over each of the eight quadrants of the bounding box, is located in the first quadrant.
@@
    Conditional-averaged quantities are computed by ensemble averaging over the $\Pi$-structures in the local frame of reference. The conditional-averaged is performed over the $\Pi$-structures for a given Reλ and $r$. Prior to performing the average, the spatial coordinates of the local frame of reference attached to each individual structure are re-scaled based on the diagonal of the rotated bounding box of the $\Pi$-structure. This process is grounded in the geometric self-similarity of the $\Pi$-structures discussed in §Geometric properties of $\Pi$-structures. The conditional-averaged $\phi$, denoted by $\{\phi\}$, is calculated as
    $\{\phi\}(\boldsymbol{\hat{x}})=\sum_{n=1}^{N}\frac{\phi\left(\boldsymbol{x}_{n}+d_{n}\boldsymbol{\hat{x}}\right)}{N},$ (3)
    where $n=1,...,N$ is the label for each $\Pi$-structure (with $N$ the total number of samples), $\boldsymbol{x}_{n}$ is the center of mass of the $n$-th $\Pi$-structure, $d_{n}$ is the diagonal length of bounding box of the $n$-th $\Pi$-structure, and $\boldsymbol{\hat{x}}$ is the space coordinate relative to the local frame of reference, re-scaled by the diagonal length.
    The conditional average approach outlined here enables the study of local-in-space flow patterns surrounding cascading events. We have avoided traditional spectral analysis due to its global-in-space nature, which renders the approach unsuitable for elucidating the structure of localized events. Note that the conditional average procedure is not merely a qualitative description of the flow, but rather a quantitative, non-trivial representation of the most probable states and patterns in the flow surrounding cascading events.
@@ Results
The dominant velocity patterns surrounding intense energy transfer events are shown in figure 6. The region of intense energy transfer is represented by a red iso-surface, corresponding to 0.5 of the maximum probability of finding a point belonging to the $\Pi$-structure, i.e., $P(\Pi>\Pi_{\text{rms}})=0.5$. The averaged shape is an elongated ellipsoid, consistent with the geometric analysis in §Geometric properties of $\Pi$-structures. The streamlines are calculated from the conditional-averaged velocity field ($\{u\},\{v\},\{w\}$) in the vicinity of the $\Pi$-structures after subtracting the mean velocity at the origin of the $\Pi$-structure. Notably, the flow exhibits a saddle point topology centered at the origin of the $\Pi$-structure, which is a characteristic signature of strain-dominated regions. Similar patterns have previously been reported in the context of sweep and ejection pairs in homogeneous shear turbulence [54] and wall-bounded turbulence [62, 63]. However, the results presented here specifically pertain to the inertial range of isotropic turbulence.
Figure 6: Streamlines of the conditional-averaged velocity surrounding intense energy cascade events. The streamlines are constrained to the planes (a) $\hat{y}=0$ and $\hat{z}=0$ (b) $\hat{x}=0$ and $\hat{z}=0$. The red isosurface is 50% probability of finding a point belonging to a $\Pi$-structure. The results are for case HIT3 and $r=30\eta$.
@@
Figure 7: (a,b) Iso-surface of the probability of finding a point belonging to a $\Pi$-structure equal to 0.5 (red) and conditional-averaged enstrophy $\{\widetilde{\omega}_{i}\}\{\widetilde{\omega}_{i}\}$ at $0.35\times\max\left(\{\widetilde{\omega}_{i}\}^{2}\right)$ at $\hat{z}>0$ (yellow) and $\hat{z}<0$ (blue) is shown in isometric view (a) and $\hat{x}$-$\hat{y}$ projection (b). The results are for case HIT3 and $r=30\eta$. (c,d) Conditional-averaged enstrophy for (c) HIT3 and $r=30\eta$ (color) and HIT2 and $r=30\eta$ (solid black line); (d) HIT3 and $r=30\eta$ (color) and HIT3 and $r=40\eta$ (solid red line). The values are normalized by the standard deviation of $\{\widetilde{\omega}_{i}\}\{\widetilde{\omega}_{i}\}$.
@@
It is important to note that the hairpin-like structure identified does not imply that instantaneous vortices are shaped in that manner. Rather, it signifies that the irregular arrays of vortices tend to be preferentially located in those regions. Two examples of the instantaneous field surrounding the $\Pi$-structures are depicted in figure 8(b). It can be seen that intense energy transfer events ($\Pi>\Pi_{\text{\text{rms}}}$, colored in red) are surrounded by areas of high enstrophy ($\widetilde{\omega}_{i}\widetilde{\omega}_{i}$, colored in green). In these particular examples, multiple vortex tubes are positioned closely to the $\Pi$-structures, yet without overlapping in space. Similar instances, consisting of regions of intense $\Pi$ surrounded by distinct regions of intense $\widetilde{\omega}_{i}\widetilde{\omega}_{i}$, were consistently observed. While certain $\Pi$-structures did overlap with vortex tubes, such occurrences were less frequent, in accordance with the results from the conditional-averaged flow.
@@
Figure 8: Examples of instantaneous intense cascade events: iso-surfaces of high transfer regions $\Pi>\Pi_{\text{rms}}$ (red) and high enstrophy regions at $0.025\times\max\left(\widetilde{\omega}_{i}\widetilde{\omega}_{i}\right)$ (green). The results are for case HIT2 and $r=30\eta$
@@
The results above rely on the assumption that only one type of flow pattern surrounds the intense energy transfer events. However, if patterns with different flow topologies are involved in the cascading process, performing an ensemble average over all the $\Pi$-structures would yield a distorted image of the coherent structure. To assess the possibility of distinct flow patterns contributing to the energy cascade, we divide the samples used to compute $\{\phi\}$ into $n$ groups with the goal of separating instances containing different flow structures. This is achieved by using proper orthogonal decomposition of the velocity field surrounding $\Pi$-structures. Then, the k-means algorithm is applied to place the samples into $n$ groups that minimize within-cluster variances of the three most energetic POD coefficients, denoted by $a_{1}$, $a_{2}$, and $a_{3}$. The POD coefficients and groups are shown in figure 9(a) for the case $n=2$. The conditional-averaged velocity field for each group is portrayed in figures 9(b) and (c). Comparison of the results indicates that both groups share similar flow features despite the division of the samples being targeted to separate velocity patterns. Although not shown, the process was repeated for $n=3$ and $n=4$ groups, which yielded similar conclusions. These findings suggest that the conditional-averaged flow structure outlined in this section is physically relevant and not merely an artifact of the averaging process. Additionally, the results in figure 9 are computed for case HIT2 and $r=30\eta$, which demonstrates that the streamlines in figure 6(a) also hold for different Reλ and filter width.
@@
$12\frac{\partial\tau_{ij}^{r}}{\partial(r^{2})}=\frac{1}{2}\nabla^{2}\tau_{ij}^{r}+\overline{A}_{ij}^{r}\overline{A}_{ij}^{r},\quad\tau_{ij}^{r=0}=0,$ (4)
where $A_{ij}=\partial u_{i}/\partial x_{j}$ is the velocity gradient tensor and $\nabla^{2}$ is the Laplacian. By solving Eq. (4) and projecting onto $\widetilde{S}_{ij}$, the kinetic energy transfer can be decomposed as

### v4

@@ Introduction
There is general agreement from the community that the answer to the first question is yes. The net transfer of energy from large to small scales is mainly accomplished by interactions among flow motions of similar size. The conclusion is supported by evidence from diverse approaches, such as scaling analysis [12, 13, 14, 15, 16, 17, 18, 19, 20], triadic interactions in Fourier space [21, 22], time correlations [23, 24], and information-theoretic causality [25, 26], to name a few.
The degree of consensus is lower regarding the second question. Several theories have been proposed to describe the physical mechanism(s) that enable the transfer of energy from larger to smaller scales throughout the inertial range. One of the first mechanisms proposed is based on the concept of vortex stretching. In this scenario, vorticity is stretched by the strain-rate either at the same scale or at a larger scale [27, 28, 29, 30, 31, 32, 33, 34]. Some authors have further proposed that the mechanistic details of the process can be explained by successive reconnections of anti-parallel vortex tubes [35, 36, 37, 38], vortex reconnection of two long, straight anti-parallel vortex tubes with localized bumps [39], and the presence of helical instabilities in vortex rings [40, 41]. These viewpoints, while not dynamically equivalent, are still compatible with the vortex-stretching driven energy cascade.
The main competing theory to the vortex stretching mechanism is the self-amplification of the rate-of-strain either by same-scale interactions or by the amplification from larger scale strain-rate [42, 43, 44, 45, 46]. In both scenarios, the strain-rate self-amplification stands as the key contributor to the transfer of energy among scales, whereas vortex stretching is merely the effect (rather than the cause) of the energy cascade. Given the kinematic relationship between vortex stretching and strain-rate self-amplification [47, 48], some authors have argued that both mechanisms play a relevant role in the dynamics of the energy cascade as one cannot occur without the other [49, 50].
@@ Methods
We use direct numerical simulations of isotropic turbulence [23]. The numerical setup corresponds to isotropically forced turbulence within a triply periodic domain at $\mathrm{Re}_{\lambda}=146,236$, and $384$, where $\mathrm{Re}_{\lambda}$ is the Reynolds number based on the Taylor microscale. The simulations are labeled as HIT1 (for $\mathrm{Re}_{\lambda}=146$), HIT2 (for $\mathrm{Re}_{\lambda}=236$), and HIT3 (for $\mathrm{Re}_{\lambda}=384$). Table 1 contains the numerical details of the simulations in terms of grid resolution, domain size, and number of snapshots. The reader is referred to Ref. [23] for additional information about the numerical setup.
@@
To investigate the structure of the energy transfer, the velocity field is decomposed into large and small scales using a low-pass Gaussian filter such that $u_{i}(\mathbf{x},t)=\widetilde{u}_{i}(\mathbf{x},t)+u^{\prime}_{i}(\mathbf{x},t)$ for $i=1,2,3$, where ${u}_{i}$ represents the $i$-th component of the instantaneous velocity, $\widetilde{u}_{i}$ is the filtered velocity, and $u^{\prime}_{i}$ corresponds to the remaining fluctuating velocity component. The spatial coordinate is $\mathbf{x}\equiv[x_{1},x_{2},x_{3}]$. Occasionally, we use the notation $[u_{1},u_{2},u_{3}]\equiv[u,v,w]$ for velocities and $[x_{1},x_{2},x_{3}]=[x,y,z]$ for space. The filter is isotropic and given in Fourier space by the Gaussian kernel $\hat{G}(\mathbf{\kappa})=\exp[-(r\kappa)^{2}/24]$, where $\kappa=\mathbf{|\kappa|}$ is the magnitude of the wavenumber vector ($\mathbf{\kappa}$), and $r$ is the filter width. Four filter widths are investigated: $r/\eta=20,30,40,$ and $60$, where $\eta=\left(\nu^{3}/\bar{\varepsilon}\right)^{1/4}$ is the Kolmogorov length-scale, $\nu$ is the kinematic viscosity, and $\bar{\epsilon}$ is the space-time averaged dissipation of turbulent kinetic energy. Figure 1 compares one instant of the unfiltered velocity and the filtered velocity for $r=30\eta$. The values of $r$ were chosen to span across the inertial range of the turbulence cascade [23]. The kinetic energy spectra $E(\kappa)$ for the three Reynolds numbers are shown in figure 2(a). As expected, the viscous and inertial range of $E(\kappa)$ collapse in Kolmogorov units, defined by $\eta$ and the Kolmogorov velocity-scale $u_{\eta}=(\nu\bar{\epsilon})^{1/4}$. The compensated kinetic energy spectra $\kappa^{5/3}E(\kappa)$ for HIT3 is included in figure 2(b) for the unfiltered case and for $r/\eta=20,30,40$ and $60$. The latter shows that the filtered widths lie within the inertial range of the moderate Reynolds numbers considered.
It is worth noting that other filtering approaches are also possible. Among them, we can mention the box filter (i.e., filter width fully localized in physical space) and the Fourier cut-off filter (i.e., filter width fully localized in wavenumber space). Although not shown, we compared the averaged energy transfer patterns using box filters and Fourier cut-off filters and obtained qualitatively similar results to those presented below. In this work, we focus our analysis on the Gaussian filter, as it offers a compromise between the box and Fourier filters while allowing for the decomposition of energy transfer mechanisms later discussed in Eq. (7).
@@
The instantaneous inter-scale kinetic energy transfer at one instance is visualized in figure 3(a). Our primary efforts are directed toward analyzing the coherent structure of the flow in the vicinity of intense forward cascade events. Regions of strong positive $\Pi$ are isolated by thresholding the field with the parameter $\alpha>0$, such that $\Pi>\alpha\Pi_{\text{rms}}$, where $\Pi_{\text{rms}}$ represents the standard deviation of $\Pi$ over all the dataset. The value of $\alpha=1$ is selected following previous work [55]. We show in the Supplementary Information that the conclusions drawn in this manuscript also hold for other values of $\alpha$. Then, individual three-dimensional $\Pi$-structures are defined as contiguous regions in space satisfying $\Pi>\Pi_{\text{rms}}$. These structures will be utilized as markers to conditionally average the flow around them. The $\Pi$-structures occupy less than 20% of the total fluid volume but contribute to 70% of the total kinetic energy transfer. Figure 3(b) visualizes the regions associated with forward (red, $\Pi>\Pi_{\text{rms}}$) and backward (blue, $\Pi<-\Pi_{\text{rms}}$) intense cascade events for a given snapshot. Note that we use $\Pi=\Pi(\mathbf{x},t)$ to designate the scalar field of inter-scale kinetic energy transfer, which is a function of both space and time, whereas the term $\Pi$-structure refers to the individual structures defined by contiguous regions in space that satisfy $\Pi(\mathbf{x},t)>\Pi_{\text{rms}}$ at a given time.
Forward cascade events dominate over their backward counterparts, as clearly appreciated in figure 3(b), consistent with previous results in the literature [58, 59, 60]. We will focus solely on the forward energy cascade, while we do not address inverse cascading events. This decision is primarily driven by two factors. Firstly, the number of intense backward cascade events is orders of magnitude smaller than the forward events, as illustrated in figure 3(b). Secondly, whereas the importance of the forward energy cascade is widely recognized in terms of dynamics and reduced-order modeling of the flow, the same consensus is not established for the inverse energy cascade in three-dimensional turbulence. Recent works have suggested that the inverse energy cascade may play a minor role in the flow dynamics or even lack physical significance [46, 56, 25, 26].
@@
We characterize the geometric properties of the $\Pi$-structures in terms of size and fractal dimensions across different filter widths and Reynolds numbers. A local frame of reference is determined using the center of mass and the principal axes of inertia for each $\Pi$-structure. The center of mass for the $n$-th $\Pi$-structure, denoted by $\mathbf{x}_{n}=[x_{n,1},x_{n,2},x_{n,3}]$, is computed assuming a solid object with constant density:
$\mathbf{x}_{n}=\frac{\iiint_{\mathcal{V}_{n}}\mathbf{x}_{p}\mathrm{d}V_{n}}{\iiint_{\mathcal{V}_{n}}\mathrm{d}V_{n}},$ (2)
where $\mathbf{x}_{p}=[x_{p,1},x_{p,2},x_{p,3}]$ represents the position vector of points within the volume of the structure, denoted by $\mathcal{V}_{n}$. The principal axes of inertia are obtained from the eigenvalues of the moment of inertia tensor for the $n$-th $\Pi$-structure:
$I^{n}_{ij}=\iiint_{\mathcal{V}_{n}}\left(\delta_{ij}l_{n}^{2}-l_{n,i}l_{n,j}\right)\mathrm{d}V_{n},$ (3)
where $l_{n,i}=x_{p,i}-x_{n,i}$ is the $i$-th component of the position vector relative to the center of mass, $l_{n}^{2}=l_{n,i}l_{n,i}$ is the squared distance from the center of mass, and $\delta_{ij}$ is the Kronecker delta.
@@
The typical shape of the $\Pi$-structures is quantified via their fractal dimension. We employed the box-counting method to calculate the Minkowski-Bouligand dimension of each $\Pi$-structure [61, 62]. For each object, the computational domain is divided into cubes with a side length of $l$, and the number $N_{b}(l)$ of boxes containing at least one point of the object is counted. In the case of a pure fractal set with dimension $D$, the number of boxes would follow a power law $N_{b}(l)\sim l^{-D}$. In practice, this relationship only holds within a very restricted range of scales, bounded by a large-scale and a small-scale cut-off. We define the local fractal dimension as
@@
The fractal dimension of the object can be taken along the range over which the slope of $D_{l}(l)$ is minimum or approximately constant.
Figure 4(d) shows the PDF of the fractal dimensions for the $\Pi$-structures for different filter widths and Reλ. Space-filling spheroidal shapes would exhibit fractal dimensions close to $D\approx 3$, whereas sheet-like or filament-like structures would have $D\approx 2$ or $D\approx 1$, respectively. The PDF of $D$ reveals a bell-shaped distribution centered around $2.45$, suggesting that the objects have an intermediate form between sheet-like and strictly spheroidal shapes. Therefore, although some $\Pi$-structures may exhibit complex shapes, most often they resemble elongated ellipsoids with arbitrary orientation. In summary, the results from this section show that $\Pi$-structures are geometrically self-similar across scales within the inertial range and for the three Reλ considered. This finding justifies the conditional averaging procedure introduced in the subsequent section.
@@
    First, individual $\Pi$-structures are identified using the thresholding approach previously discussed. Examples of three individual $\Pi$-structures are shown in figure 5. A characteristic length scale is assigned to each $\Pi$-structure, defined by the diagonal of their bounding box, $\sqrt{l_{x}^{2}+l_{y}^{2}+l_{z}^{2}}$. Our interest lies in studying structures that are of a size similar to the filter width. $\Pi$-structures larger than half the length of the entire computational domain are excluded from the analysis. Similarly, structures smaller than $3^{3}$ grid points are also disregarded. The number of valid $\Pi$-structures considered for the analysis is roughly on the order of $\mathcal{O}(10^{4})$ for each Reλ and filter width. Given the large number of samples, the statistical uncertainty of the results were found to be small and more details are discussed in the Supplementary Information.
@@
    Each individual $\Pi$-structure is assigned a local frame of reference, with the origin positioned at its center of mass. This is the same frame of reference used in the previous section to compute the characteristic lengths of the $\Pi$-structures. The axes directions are aligned with the principal axes of inertia of the structure relative to the center of mass. A rotation matrix, constructed from these eigenvectors, is used to rotate the flow field surrounding each structure. Figure 5 illustrates the rotated bounding boxes for three $\Pi$-structures. The principal axes of inertia introduced above does not inherently define the positive or negative directions of the axes. To resolve this ambiguity, we set the direction of the axes such that the highest enstrophy $\widetilde{\omega}_{i}\widetilde{\omega}_{i}$, averaged over each of the eight quadrants of the bounding box, is located in the first quadrant.
@@
    Conditional-averaged quantities are computed by ensemble averaging over the $\Pi$-structures in the local frame of reference. The conditional-averaged is performed over the $\Pi$-structures for a given Reλ and $r$. Prior to performing the average, the spatial coordinates of the local frame of reference attached to each individual structure are re-scaled based on the diagonal of the rotated bounding box of the $\Pi$-structure. This process is grounded in the geometric self-similarity of the $\Pi$-structures discussed in above. The conditional-averaged $\phi$, denoted by $\{\phi\}$, is calculated as
    $\{\phi\}(\mathbf{\hat{x}})=\sum_{n=1}^{N}\frac{\phi\left(\mathbf{x}_{n}+d_{n}\mathbf{\hat{x}}\right)}{N},$ (5)
    where $n=1,...,N$ is the label for each $\Pi$-structure (with $N$ the total number of samples), $\mathbf{x}_{n}$ is the center of mass of the $n$-th $\Pi$-structure, $d_{n}$ is the diagonal length of bounding box of the $n$-th $\Pi$-structure, and $\mathbf{\hat{x}}$ is the space coordinate relative to the local frame of reference, re-scaled by the diagonal length.
    The conditional average approach outlined here enables the study of local-in-space flow patterns surrounding cascading events. We have avoided traditional spectral analysis due to its global-in-space nature, which renders the approach unsuitable for elucidating the structure of localized events. Note that the conditional average procedure is not merely a qualitative description of the flow, but rather a quantitative representation of the most probable states and patterns in the flow surrounding cascading events.
@@ Results
The dominant velocity patterns surrounding intense energy transfer events are shown in figure 6. The region of intense energy transfer is represented by a red iso-surface, corresponding to 0.5 of the maximum probability of finding a point belonging to the $\Pi$-structure, i.e., $P(\Pi>\Pi_{\text{rms}})=0.5$. The averaged shape is an elongated ellipsoid, consistent with our geometric analysis. The streamlines are calculated from the conditional-averaged velocity field ($\{\tilde{u}\},\{\tilde{v}\},\{\tilde{w}\}$) in the vicinity of the $\Pi$-structures after subtracting the mean velocity at the origin of the $\Pi$-structure. Notably, the flow exhibits a saddle point topology centered at the origin of the $\Pi$-structure, which is a characteristic signature of strain-dominated regions. Similar patterns have previously been reported in the context of sweep and ejection pairs in homogeneous shear turbulence [55] and wall-bounded turbulence [63, 64]. However, the results presented here specifically pertain to the inertial range of isotropic turbulence.
Figure 6: Streamlines of the conditional-averaged velocity ($\{\tilde{u}\},\{\tilde{v}\},\{\tilde{w}\}$) surrounding intense energy cascade events. The streamlines are constrained to the planes (a) $\hat{y}=0$ and $\hat{z}=0$ (b) $\hat{x}=0$ and $\hat{z}=0$. The red isosurface is 50% probability of finding a point belonging to a $\Pi$-structure. The results are for case HIT3 and $r=30\eta$.
@@
Figure 7: (a,b) Iso-surface of the probability of finding a point belonging to a $\Pi$-structure equal to 0.5 (red) and conditional-averaged enstrophy $\{\widetilde{\omega}_{i}\}\{\widetilde{\omega}_{i}\}$ at $0.35\times\max\left(\{\widetilde{\omega}_{i}\}^{2}\right)$ at $\hat{z}>0$ (yellow) and $\hat{z}<0$ (blue). The results are shown in isometric view in panel (a) and in the $\hat{x}$-$\hat{y}$ view in panel (b). The results are for case HIT3 and $r=30\eta$. (c,d) Conditional-averaged enstrophy for (c) HIT3 and $r=30\eta$ (color) and HIT2 and $r=30\eta$ (solid black line); (d) HIT3 and $r=30\eta$ (color) and HIT3 and $r=40\eta$ (solid red line). The values are normalized by the standard deviation of $\{\widetilde{\omega}_{i}\}\{\widetilde{\omega}_{i}\}$.
@@
It is important to note that the hairpin-like structure identified does not imply that instantaneous vortices are shaped in that manner. Rather, it signifies that the irregular arrays of vortices tend to be preferentially located in those regions. Two examples of the instantaneous field surrounding the $\Pi$-structures are depicted in figure 8. It can be seen that intense energy transfer events ($\Pi>\Pi_{\text{\text{rms}}}$, colored in red) are surrounded by areas of high enstrophy ($\widetilde{\omega}_{i}\widetilde{\omega}_{i}$, colored in green). In these particular examples, multiple vortex tubes are positioned closely to the $\Pi$-structures, yet without overlapping in space. Similar instances, consisting of regions of intense $\Pi$ surrounded by distinct regions of intense $\widetilde{\omega}_{i}\widetilde{\omega}_{i}$, were consistently observed. While certain $\Pi$-structures did overlap with vortex tubes, such occurrences were less frequent, in accordance with the results from the conditional-averaged flow.
@@
Figure 8: Examples of instantaneous intense cascade events: iso-surfaces of high transfer regions $\Pi>\Pi_{\text{rms}}$ (red) and high enstrophy regions at $0.025\times\max\left(\widetilde{\omega}_{i}\widetilde{\omega}_{i}\right)$ (green). The results are for case HIT2 and $r=30\eta$.
@@
The results above rely on the assumption that only one type of flow pattern surrounds the intense energy transfer events. However, if patterns with different flow topologies are involved in the cascading process, performing an ensemble average over all the $\Pi$-structures would yield a distorted image of the coherent structure. To assess the possibility of distinct flow patterns contributing to the energy cascade, we divide the samples used to compute $\{\phi\}$ into $n$ groups with the goal of separating instances containing different flow structures. The goal is to create groups of samples that are as distinct as possible. This is accomplished by first extracting the large-scale velocity patterns surrounding the $\Pi$-structures using proper orthogonal decomposition (POD), followed by applying the k-means algorithm to maximize the differences between samples in different groups. The POD modes are computed for the velocity vector $[\tilde{u},\tilde{v},\tilde{w}]$ after subtracting the mean velocity at the origin of the $\Pi$-structure. The spatial extent of the domain for computing the POD modes is defined by a cube with a side length of 1 and origin at the center of mass of the structure. As shown in figure 7, this domain is sufficient to enclose the key flow patterns. The first three most energetic POD modes are used to characterize the large-scale flow pattern for each $\Pi$-structure. The corresponding POD coefficients are denoted as $a_{1}(n)$, $a_{2}(n)$, and $a_{3}(n)$, where $n$ refers to the sample (i.e., the $\Pi$-structure) they belong to. The k-means algorithm is applied to place the samples into $n$ groups that minimize within-cluster variances of the three most energetic POD coefficients. The POD coefficients and groups are shown in figure 9(a) for the case $n=2$. The conditional-averaged velocity field for each group is portrayed in figures 9(b) and (c). Comparison of the results indicates that both groups share similar flow features despite the division of the samples being targeted to separate velocity patterns. Although not shown, the process was repeated for $n=3$ and $n=4$ groups, which yielded similar conclusions. These findings suggest that the conditional-averaged flow structure outlined in this section is physically relevant and not merely an artifact of the averaging process. Additionally, the results in figure 9 are computed for case HIT2 and $r=30\eta$, which demonstrates that the streamlines in figure 6(a) also hold for different Reλ and filter widths.
@@
$12\frac{\partial\tau_{ij}^{r}}{\partial(r^{2})}=\frac{1}{2}\nabla^{2}\tau_{ij}^{r}+\overline{A}_{ik}^{r}\overline{A}_{jk}^{r},\quad\tau_{ij}^{r=0}=0,$ (6)
where $A_{ij}=\partial u_{i}/\partial x_{j}$ is the velocity gradient tensor and $\nabla^{2}$ is the Laplacian operator. By solving Eq. (6) and projecting onto $\widetilde{S}_{ij}$, the kinetic energy transfer can be decomposed as